OpenStudy (anonymous):

WILL AWARD MEDAL! Can anyone solve this? 675.54=P(1+0.025/2)^2(11)?

OpenStudy (anonymous):

Like this? \[675.54=p(1+0.025/2)^{2(11)}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Exactly like this?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

You need to solve for p I assume?

OpenStudy (anonymous):

i believe so

OpenStudy (anonymous):

Well we need to isolate p to one side of the equation. Any thoughts on how?

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

8 = 4x. If I divide both sides by 4 what do I end up with?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

2=x correct

OpenStudy (anonymous):

So if I divide both sides of you equation by the following..what do you end up with?\[(1+0.025/2)^{2(11)}\]

OpenStudy (anonymous):

(1+0.025/2)^2(11) (1.0125)^2(11)?

OpenStudy (anonymous):

(1.0125)^22?

OpenStudy (anonymous):

That's one step ahead of me. I f we divide both sides of the equation y (1+0.025/2)2^(11)..we would end up with \[\frac{ 675 }{ (1+0.025/2)^{2(11)} } = p\]

OpenStudy (anonymous):

Do you see how I did that?

OpenStudy (anonymous):

okay then what?

OpenStudy (anonymous):

Solve the equation. First solve for (1.0125) ^22....then divide that into 675 and there is your answer for p.

OpenStudy (anonymous):

this is what i get for (1.0125)^22=1.3142884808187545

OpenStudy (anonymous):

@nelsonjedi ?

OpenStudy (anonymous):

675.54=11P(1.0125)^2 I think...